Circular No. 9064

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V2672 OPHIUCHI = NOVA OPHIUCHI 2009 S. Nakano, Sumoto, Japan; and H. Yamaoka, Kyushu University, re-

port the discovery by K. Itagaki (Teppo-cho, Yamagata, Japan) of an apparent nova (mag 10.0) on unfiltered CCD survey frames taken on Aug. 16.515 UT using a 0.21-m reflector (limiting mag 15), the position given as α = 17^{h}38^{m}19^{s}.68, δ = −26^{o}44′14′′.0 (equinox 2000.0). Additional unfiltered magnitudes and position end figures: 2008 June 10.645, [15.0 (Itagaki); 2009 Aug. 16.526, 10.1, 19^{s}.70, 14′′.0 (Itagaki, 0.60-m reflector); 16.576, 10.2, 19^{s}.70, 13′′.8 (K. Kadota, Ageo, Japan, 0.25-m reflector); 17.428, ∼ 11.1 (Itagaki). Both Itagaki and Kadota find nothing at this position on the Digitized Sky Survey (Kadota citing red DSS images from 1991 and 1996). Yamaoka finds no object brighter than mag about 14.0 on a ASAS-3 V -band image taken on 2009 Aug. 14.142. Other observers have reported additional measured positions and magnitudes, as reported on CBET 1910, including M. Nissinen and V.-P. Hentunen (Varkaus, Finland), who could find nothing visible at the position of the nova on a Palomar Sky Survey DSS image (limiting red mag 20.8). Visual magnitude estimate by P. Schmeer, Bischmisheim, Germany: Aug. 17.854, 12.6.

Low-resolution spectroscopy (range 450–800 nm) by K. Ayani, N. Murakami, K. Hata, A. Tanaka, M. Tachibana, and A. Kanda with the Bisei Astronomical Observatory 1.01-m telescope on Aug. 17.6 UT (details on CBET 1911) reveals broad, prominent Hα emission (FWHM ∼ 8000 km/s; equivalent width ∼ 49 nm) upon a red continuum; broad Hβ emission line also was present. Medium-resolution spectroscopy (range 540–670 nm) obtained by U. Munari, T. Saguner, P. Ochner, A. Siviero, A. Maitan, P. Valisa, S. Dallaporta, and S. Moretti on Aug. 17.83 with the Asiago Astrophysical Observatory 1.22-m telescope (details on CBET 1912) is dominated by a very strong and highly structured Hα in emission (velocity width 11500 km/s at the base) on a featureless continuum with interstellar Na I and diffuse intertellar bands (at 577.9, 628.4 and 661.4 nm) in strong absorption. An low-resolution CCD spectrogram (range 400–870 nm) obtained by Munari et al. on Aug. 17.87 with a 0.6-m telescope at Varese show Hα, Hβ, O I 844.6-nm, and possibly He I 706.5-nm emission lines. Their photometry shows that the nova faded from V = 12.63 on Aug. 17.47 to V = 13.05 on Aug. 18.49; they add that this appears to be a highly reddened outburst occurring on a massive white dwarf, not dissimilar from the U-Sco type of recurrent novae, noting that a search of plate archives for missed previous outbursts could be useful. N. Samus and E. V. Kazarovets report that the GCVS team has assigned the designation V2672 Oph to this nova.

2009 August 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green